OpenStudy (keana):

Which ratio forms a proportion with 18/38? a: 9/29 b: 7/27 c: 24/44 d: 27/57

Parth (parthkohli):

Look at 18/38 as the fraction \(\Large {18 \over 38}\). Now find its equivalent fraction. Do you know how to?

OpenStudy (anonymous):

ans is 9/19

OpenStudy (anonymous):

Is there any problem in solving that??? 18 = 2*9 38 = 2*19 And then use your little bit of brain and then solve it..

OpenStudy (keana):

it cannot be 9/19

Parth (parthkohli):

For example check: \( \color{Black}{\Rightarrow 18:38=27:37?}\)

OpenStudy (anonymous):

@arch1 why you are telling answers directly??

OpenStudy (anonymous):

Please read the question before giving out the wrong answer!

OpenStudy (keana):

it can't be 9/19 !

OpenStudy (anonymous):

@ParthKohli there it will come 57 in place of 37..

OpenStudy (anonymous):

Step 1: Reduce the fraction (if possible) to the smallest form. Step 2: Multiply by a form of 1, e.g. 1/1, 2/2, etc. The results have the same proportion (or ratio).

Parth (parthkohli):

lol I was correcting that ;P

OpenStudy (anonymous):

oops ans is 27/57

Parth (parthkohli):

Oh it is 57

Parth (parthkohli):

\( \color{Black}{\Rightarrow 27:57 = 9:19}\)

OpenStudy (anonymous):

Please stop giving out the answer.

OpenStudy (anonymous):